# Tensor products of bimodules and bimodule pairs over monoids

### v2

@@ 4.1.1. Tensor products of prepairs
$((\mathcal{M}_{1}\otimes\mathcal{M}_{2}),(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-prepair in (i), and a pair in (ii).
@@
$((\mathcal{M}_{1}\otimes\mathcal{M}_{2}),(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0})$ is closed under left multiplication by $\mathcal{T}_{1}$ and right multiplication by $\mathcal{T}_{2}.$ $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}$ is closed under addition in (ii). ∎
@@
We have the $\mathcal{T}_{1}\times_{\mathcal{T}}\mathcal{T}_{2}$ action, which induces a $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$-action since $(a_{1}a\otimes a_{2})b=a_{1}aba_{2}=a_{1}baa_{2}=(a_{1}\otimes aa_{2})b.$
@@ 4.2. Properties of tensor products
When $\mathcal{M}_{1},\mathcal{M}_{2}$ have negation maps, then a balanced map is required to satisfy $((-)v)\otimes w=v\otimes(-w),$ which we define to be $(-)(v\otimes w)$.
Proposition 4.13.
@@
given by $(f_{1}\otimes f_{2})(v\otimes w)=f_{1}(v)\otimes f_{2}(w).$
@@
The balanced map $\Psi:(v,w)\mapsto f_{1}(v)\otimes f_{2}(w)$ sends $\Phi_{1}$ to $\Phi_{2},$ seen by applying $f_{1}$ and $f_{2}$ to the respective components. ∎
@@
    If $\mathcal{M}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a $\mathcal{T}_{1}$-module, with the $\mathcal{M}_{1}$-action $a(v\otimes w)=av\,\otimes w$. Analogously for $\mathcal{M}_{2}.$
@@
      $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule, with the actions $a_{1}(v\otimes w)a_{2}=a_{1}v\,\otimes\,wa_{2}$.
@@
    If $\mathcal{T}$ is central in both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, then $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\cong\mathcal{M}_{2}\otimes_{\mathcal{T}}\mathcal{M}_{1}$ via $(v_{1}\otimes v_{2})\mapsto v_{2}\otimes v_{1}.$
@@
Repeated applications of Proposition 4.13.
@@
(iii) $f$ is determined by the balanced map $\psi:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}$ given by $\psi(v,w)=f(v\otimes w).$
(iv) Take the balanced map $(v_{1},v_{2})\mapsto(v_{2}\otimes v_{1})$.
@@
    Associativity of addition in $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ follows from associativity in $\mathcal{F}_{1}\times\mathcal{F}_{2}.$
@@
The tensor product in this case satisfies the criteria of [21, Definitions 1,2].
@@
We shall show that $S\otimes(S_{1}^{\prime}+S_{2}^{\prime})\subseteq SS_{1}^{\prime}\,+\,SS_{2}^{\prime}$ for all $S\subseteq\mathcal{H}_{1}$ and $S_{1}^{\prime},S_{2}^{\prime}\subseteq\mathcal{H}_{2}$. By definition, $S\otimes(S_{1}^{\prime}+S_{2}^{\prime})=\{a\otimes a^{\prime}:a\in S,\ a^{\prime}\in S_{1}^{\prime}+S_{2}^{\prime}\}=\{a(S_{1}^{\prime}+S_{2}^{\prime}):a\in S\}.$ Thus we need to show that $a\otimes(S_{1}^{\prime}+S_{2}^{\prime})\subseteq a\otimes S_{1}^{\prime}\,+\,a\otimes S_{2}^{\prime}$ for each $a\in S.$ But for each $a^{\prime}\in S_{1}^{\prime}+S_{2}^{\prime}$ if $a^{\prime}\in b_{1}^{\prime}+b_{2}^{\prime}$ for suitable nonzero $b_{i}^{\prime}\in S_{2}^{\prime}$, then $a\otimes a^{\prime}=(a\otimes b_{1}^{\prime})+(a\otimes b_{2}^{\prime}),$ and we are done by induction on height. Hence we may assume that $a^{\prime}\in\mathcal{H}_{2},$ and the assertion is immediate. ∎
4.3. Functorial properties of residue tensors
The notation for hyperaddition in residue hypermodules continues to be $\boxplus$.⊞ .
Lemma 4.21.
Suppose that $\mathcal{T}_{1},\mathcal{T}_{2}$ are abelian monoids, $\mathcal{M}_{i}$ are $\mathcal{T}_{i}$-modules for $i=1,2,$ and $G$ is a common subgroup of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}.$ The tensor product of the quotient hypermodules $\mathcal{H}_{1}=\mathcal{M}_{1}/G_{1}$ and $\mathcal{M}_{2}=\mathcal{M}_{2}/G_{2}$ is isomorphic to $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})/(G_{1}\otimes G_{2})$.
Proof.
4.4. Tensor extensions of a pair
There are two flavors to extending the base monoid $\mathcal{T}$.
@@ 4.4. Tensor extensions of a pair
  1. (i)
    Define the $\mathcal{T}$-tensor extension $\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{M}$ to be the semigroup $(\mathcal{\mathcal{T}}^{\prime}\times F(\mathcal{M})/\Phi,$ where $\Phi$ is the congruence generated by all
    $\bigg{(}\big{(}a^{\prime},v+w\big{)},\bigg{(}(a^{\prime}a,w),(a^{\prime},aw)\bigg{)}$ (4.9)
    $\forall v,w\in\mathcal{\mathcal{M}}_{i},$ $a^{\prime}\in\mathcal{T}$.
  2. (ii)
    When $\mathcal{M}$ is left (weakly) admissible, define the left (weakly) admissible $\mathcal{T}$-tensor extension $\mathcal{T}^{\prime}\otimes_{\mathcal{T};\operatorname{adm}}\mathcal{M}$ to be the semigroup $(\mathcal{F}(\mathcal{T}^{\prime}\times\mathcal{M})/\Phi,$ where $\Phi$ is the congruence generated by all
    $\bigg{(}\big{(}a^{\prime},v+w\big{)},(a^{\prime},v)+(a^{\prime},w)\bigg{)},\qquad\bigg{(}(a^{\prime}a,w),(a^{\prime},aw)\bigg{)},\qquad\forall v,w\in\mathcal{\mathcal{M}}_{i},\ a^{\prime}\in\mathcal{T}.$ (4.10)
  3. (iii)
    When $\mathcal{M}$ has a negation map, we incorporate the negation map into the definition of tensor product, by enlarging $\Phi$ to include also $(\big{(}(-)a,w),\big{(}a,(-)w)\big{)}$ for $a\in\mathcal{T},$ $w\in\mathcal{F}_{i}.$
@@
    A balanced map is a function $\Psi:\mathcal{T}\times\mathcal{M}\to\mathcal{N}$ where $\mathcal{N}$ is a semigroup, satisfying
    $\Psi(a,w_{1}+w_{2})=\Psi(a,w_{1})+\Psi(a,w_{2}),$ (4.11)
@@
    for all $a\in\mathcal{T},$ $v_{i},w_{i}\in\mathcal{M}_{i}.$
@@
The map $\mathcal{T}^{\prime}\times\mathcal{M}^{\prime}\to\mathcal{M}^{\prime}$ given by $(a,\sum a_{i}\otimes y_{i})\mapsto\sum aa_{i}\otimes y_{i}$ factors through the congruence $\Phi.$
As in Proposition 4.13, we have the key observation leading to the isomorphism theorems.
@@
    Suppose $\mathcal{T}_{1},\mathcal{T}_{2}$ are monoids containing $\mathcal{T}$, and $\mathcal{M}_{i}$, are $\mathcal{T}$-modules for $i=1,2$. If $f_{1}:\mathcal{T}_{1}\to\mathcal{T}_{2}$ is a monoid homomorphism fixing $\mathcal{T}$ and $f_{2}:\mathcal{M}_{2}\to\mathcal{N}_{2}$ is a module homomorphism, then there is a $\mathcal{T}$-module homomorphism
@@
    $f_{1}\otimes f_{2}:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{1}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{N}_{2}$
@@
(ii) Consider the balanced map $f_{a_{2};w_{2}}\mathcal{T}_{1}\times\mathcal{M}\to\mathcal{T}_{1}\otimes\mathcal{M}$ given by $(a,w)\mapsto(aa_{2})\otimes(aw_{2})$. This induces a map $\mathcal{T}_{1}\otimes\mathcal{M}\to\mathcal{T}_{1}\otimes\mathcal{M}$.
@@
$1\otimes f:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{1}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{M}_{2}$
@@
    If $\mathcal{T}\subseteq\mathcal{T}_{1}\subseteq\mathcal{T}_{2}$ then $\mathcal{T}_{2}\otimes(\mathcal{T}_{1}\otimes\mathcal{A})\cong\mathcal{T}_{2}\otimes\mathcal{A}.$
@@
    If $\mathcal{A}_{2}\to\mathcal{A}_{1}$ is an epimorphism of $\mathcal{T}$-semialgebras, then $\mathcal{T}_{1}\otimes\mathcal{A}_{2}\to\mathcal{T}_{1}\otimes\mathcal{A}_{1}$ is an epimorphism.
@@ 5. Tensor products of morphisms which need not be homomorphisms
The proof of Proposition 4.13 fails for morphisms which are not homomorphisms, and thus its usefulness for hypermagmas is limited. Our goal nevertheless is to salvage whatever information that we can. We do not even have a functor for residue hypermagmas.
@@
Suppose $f_{i}\mathcal{M}_{i}\to\mathcal{N}_{i}$ are homomorphisms of $\mathcal{T}_{i}$-modules, for $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ abelian, and $G$ is a common subgroup of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}.$ Then writing $\overline{\mathcal{M}_{i}}=\mathcal{M}_{i}/G$, $\overline{\mathcal{N}_{i}}=\mathcal{N}_{i}/G,$ and $\bar{f}_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}/G$ for the weak morphism induced by $f_{i}$, we would want a balanced map $\Psi:=\bar{f}_{1}\times\bar{f}_{2}:{\mathcal{M}_{1}}\times{\mathcal{M}_{2}}\to\overline{\mathcal{N}_{1}}\otimes\overline{\mathcal{N}_{2}}$ given by
@@
But $\Psi(v_{1}+w_{1},x_{2})=f_{1}(v_{1}+w_{1})Gf_{2}(x_{2})$ whereas $\Psi(v_{1},x_{2})\boxplus\Psi(w_{1},x_{2})=f_{1}(v_{1})Gf_{2}(x_{2})\boxplus f_{1}(w_{1})Gf_{2}(x_{2}),$ which could be larger.
@@
In [20], to enable a categorical approach for hypermagmas, the tensor product of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ is defined as satisfying
@@
If $v_{1}+v_{2}=v_{1}$ and $w_{1}+w_{2}=w_{1}$ then $((v_{1}\otimes w_{1})+(v_{1}\otimes w_{2}))+((v_{2}\otimes w_{1})+(v_{2}\otimes w_{2}))=(v_{1}\otimes w_{1})+(v_{2}\otimes w_{1})=(v_{1}\otimes w_{1})$ whereas
$(v_{1}\otimes w_{1})+(((v_{1}\otimes w_{2}))+(v_{2}\otimes w_{1}))+(v_{2}\otimes w_{2}))=(v_{1}\otimes w_{1})+(\emptyset+(v_{2}\otimes w_{2}))=\emptyset.$
Thus, although [20] is the appropriate approach for $\preceq$-morphisms in the category of hypermagmas, we feel that the more traditional approach may be needed in the category of hypermodules.

### v4

@@ 4.1.1. Tensor products of bimodule prepairs
$((\mathcal{M}_{1}\otimes\mathcal{M}_{2}),(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0})$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule prepair.
@@
$((\mathcal{M}_{1}\otimes\mathcal{M}_{2}),(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0})$ is closed under left multiplication by $\mathcal{T}_{1}$ and right multiplication by $\mathcal{T}_{2}.$ $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}$ is closed under addition. ∎
@@
We have the $\mathcal{T}_{1}\times\mathcal{T}_{2}$ action, which induces a $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$-action since $(a_{1}a\otimes a_{2})b=a_{1}aba_{2}=a_{1}baa_{2}=(a_{1}\otimes aa_{2})b.$
@@
Definition 4.8.
Suppose $(\mathcal{M}_{1},{\mathcal{M}_{1}}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule prepair and $(\mathcal{M}_{2},{\mathcal{M}_{2}}_{0})$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule prepair. Define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}:=({\mathcal{M}_{1}}_{0}\otimes_{\mathcal{T}}\mathcal{M}_{2})+(\mathcal{M}_{1}\otimes_{\mathcal{T}}{\mathcal{M}_{2}}_{0}).$
@@ 4.2. Properties of tensor products
When $\mathcal{M}_{1},\mathcal{M}_{2}$ are given with negation maps, then a balanced map is required to satisfy $((-)v_{1})\otimes v_{2}=v_{1}\otimes(-v_{2}),$ which we define to be $(-)(v_{1}\otimes v_{2})$.
Theorem 4.12.
@@
given by $(f_{1}\otimes f_{2})(v_{1}\otimes v_{2})=f_{1}(v_{1})\otimes f_{2}(v_{2}).$
@@
The balanced map $\Psi:(v_{1},v_{2})\mapsto f_{1}(v_{1})\otimes f_{2}(v_{2})$ sends $\Phi_{1}$ to $\Phi_{2},$ seen by applying $f_{1}$ and $f_{2}$ to the respective components. ∎
@@
    If $\mathcal{M}_{1}$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule, then $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a $\mathcal{T}_{1}$-module, with the $\mathcal{M}_{1}$-action $a(v_{1}\otimes v_{2})=av_{1}\,\otimes v_{2}$. Analogously for $\mathcal{M}_{2}.$
@@
      $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is a $(\mathcal{T}_{1},\mathcal{T}_{2})$-bimodule, with the actions $a_{1}(v_{1}\otimes v_{2})a_{2}=a_{1}v_{1}\,\otimes\,v_{2}a_{2}$.
@@
    If both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $\mathcal{T}$-bimodules, then $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}\cong\mathcal{M}_{2}\otimes_{\mathcal{T}}\mathcal{M}_{1}$ via $(v_{1}\otimes v_{2})\mapsto v_{2}\otimes v_{1}.$
@@
Repeated applications of Theorem 4.12.
@@
(iii) $f$ is determined by the balanced map $\psi:\mathcal{M}_{1}\times\mathcal{M}_{2}\to\mathcal{N}$ given by $\psi(v_{1},v_{2})=f(v_{1}\otimes v_{2}).$
(iv) Take the balanced map $(v_{1},v_{2})\mapsto v_{2}\otimes v_{1}$.
@@
    Associativity of addition in $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ follows from associativity of addition in $\mathcal{S}(\mathcal{M}_{1})\times\mathcal{S}(\mathcal{M}_{2}).$
@@
The tensor product in this subsection satisfies the criteria of [21, Definitions 1,2].
@@
We shall show that $S\otimes(S_{1}^{\prime}\boxplus S_{2}^{\prime})\subseteq SS_{1}^{\prime}\,\boxplus\,SS_{2}^{\prime}$ for all $S\subseteq\mathcal{H}_{1}$ and $S_{1}^{\prime},S_{2}^{\prime}\subseteq\mathcal{H}_{2}$. By definition, $S\otimes(S_{1}^{\prime}\boxplus S_{2}^{\prime})=\{a\otimes a^{\prime}:a\in S,\ a^{\prime}\in S_{1}^{\prime}\boxplus S_{2}^{\prime}\}=\{a(S_{1}^{\prime}\boxplus S_{2}^{\prime}):a\in S\}.$ Thus we need to show that $a\otimes(S_{1}^{\prime}\boxplus S_{2}^{\prime})\subseteq a\otimes S_{1}^{\prime}\,\boxplus\,a\otimes S_{2}^{\prime}$ for each $a\in S.$ But for each $a^{\prime}\in S_{1}^{\prime}\boxplus S_{2}^{\prime}$, if $a^{\prime}\in b_{1}^{\prime}\boxplus b_{2}^{\prime}$ for suitable nonzero $b_{i}^{\prime}\in S_{2}^{\prime}$, then $a\otimes a^{\prime}\in(a\otimes b_{1}^{\prime})\boxplus(a\otimes b_{2}^{\prime}),$ and we are done by induction on height. Hence we may assume that $a^{\prime}\in\mathcal{H}_{2},$ and the assertion is immediate. ∎
4.3. Tensor extensions
@@
  1. (1)
    Define the $\mathcal{T}$-tensor extension $\mathcal{T}^{\prime}\otimes_{\mathcal{T}}\mathcal{M}$ to be the semigroup $\mathcal{S}(\mathcal{T}^{\prime}\times\mathcal{M})/\Phi,$ where $\Phi$ is the congruence generated by all
    $\bigg(\big(a^{\prime},v+w\big),\big(a^{\prime},v\big)+\big(a^{\prime},w\big)\bigg)\qquad\bigg((a^{\prime}a,w),(a^{\prime},aw)\bigg)\qquad\forall a,a^{\prime}\in\mathcal{T},\ v,w\in\mathcal{\mathcal{M}}.$ (4.7)
    When $\mathcal{M}$ has a negation map, we incorporate the negation map into the definition of tensor product, by enlarging $\Phi$ to include also $(\big((-)a,w),\big(a,(-)w)\big)$ for $a\in\mathcal{T},$ $w\in\mathcal{M}.$
@@
    A balanced map is a function $\Psi:\mathcal{T}^{\prime}\times\mathcal{M}\to\mathcal{N}$ where $\mathcal{N}$ is a semigroup, satisfying
    $\Psi(a^{\prime},w_{1}+w_{2})=\Psi(a^{\prime},w_{1})+\Psi(a^{\prime},w_{2}),$ (4.8)
@@
    for all $a\in\mathcal{T},$ $a^{\prime}\in\mathcal{T}^{\prime},$ $v_{i},w_{i}\in\mathcal{M}_{i}.$
@@
The map $\mathcal{T}^{\prime}\times\mathcal{M}\to\mathcal{M}^{\prime}$ given by $(a,\sum a_{i}\otimes y_{i})\mapsto\sum aa_{i}\otimes y_{i}$ factors through the congruence $\Phi.$
As in Theorem 4.12, we have the key observation leading to the isomorphism theorems.
@@
    Suppose $\mathcal{T}_{1},\mathcal{T}_{2}$ are monoids containing $\mathcal{T}$, and $\mathcal{M}_{i}$ are $\mathcal{T}$-modules for $i=1,2$. If $f_{1}:\mathcal{T}_{1}\to\mathcal{T}_{2}$ is a monoid homomorphism fixing $\mathcal{T}$ and $f_{2}:\mathcal{M}_{1}\to\mathcal{M}_{2}$ is a module homomorphism, then there is a $\mathcal{T}$-module homomorphism
@@
    $f_{1}\otimes f_{2}:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{1}\to\mathcal{T}_{2}\otimes_{\mathcal{T}}\mathcal{A}_{2}$
@@
  4. (iv)
    If $\mathcal{T}_{1}$ is the disjoint union $\underset{\cdot}{\cup}\,c_{i}\mathcal{T}$ then every element of $\mathcal{T}_{1}\otimes\mathcal{M}_{2}$ can be written in the form $\sum_{i\in I}a_{i}\otimes b_{i}$, for uniquely determined $b_{i}\in\mathcal{M}_{2}$.
@@
(ii) The projection $\pi_{j}:\mathcal{T}_{1}\to\mathcal{T}$ onto the coefficient of $c_{j}$ yields a projection $1\otimes\pi_{j}:\mathcal{T}_{1}\otimes\mathcal{M}_{2}\to\mathcal{M}_{2}$. If $\sum_{i\in I}a_{i}\otimes b_{i}=\sum_{i\in I}a_{i}\otimes b^{\prime}_{i}$ then applying $\pi_{j}\otimes 1$ yields $b_{j}=b^{\prime}_{j}.$
@@
(iv) Consider the balanced map $f_{a_{2};w_{2}}\mathcal{T}_{1}\times\mathcal{M}\to\mathcal{T}_{1}\otimes\mathcal{M}$ given by $(a,w)\mapsto(aa_{2})\otimes(aw_{2})$. This induces a map $\mathcal{T}_{1}\otimes\mathcal{M}\to\mathcal{T}_{1}\otimes\mathcal{M}$. ∎
@@
$1\otimes f:\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{1}\to\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$
@@
    If $\mathcal{T}\subseteq\mathcal{T}_{1}\subseteq\mathcal{T}_{2},$ then $\mathcal{T}_{2}\otimes(\mathcal{T}_{1}\otimes\mathcal{A})\cong\mathcal{T}_{2}\otimes\mathcal{A}.$
@@
    If $\mathcal{A}_{1}\to\mathcal{A}_{2}$ is an epimorphism of $\mathcal{T}$-semialgebras, then $\mathcal{T}_{1}\otimes\mathcal{A}_{1}\to\mathcal{T}_{1}\otimes\mathcal{A}_{2}$ is an epimorphism.
@@
$4.$ Residue tensors viewed functorially
One nice feature of this approach is that the residue functor preserves tensor products. The notation for hyperaddition in residue hypermodules continues to be $\boxplus.$
Proposition 4.27.
Suppose that $\mathcal{T}_{1},\mathcal{T}_{2}$ are abelian monoids, $\mathcal{M}_{i}$ are $\mathcal{T}_{i}$-modules for $i=1,2,$ and $G$ is a common subgroup of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}.$ The tensor product of the residue hypermodules $\mathcal{H}_{1}=\mathcal{M}_{1}/G_{1}$ and $\mathcal{M}_{2}=\mathcal{M}_{2}/G_{2}$ is isomorphic to $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})/(G_{1}\otimes G_{2})$.
Proof.
@@ 5. Tensor products of morphisms which need not be homomorphisms
Now we get to the sticky part. The proof of Theorem 4.12 fails for morphisms which are not homomorphisms, and thus its usefulness for hypersemigroups is limited. Our goal nevertheless is to salvage whatever information that we can. We do not even have an obvious functor for residue hypersemigroups in general.
@@
Suppose $f_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}$ are homomorphisms of $\mathcal{T}_{i}$-modules, for $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ abelian, and $G$ is a common subgroup of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}.$ Then writing $\overline{\mathcal{M}_{i}}=\mathcal{M}_{i}/G$, $\overline{\mathcal{N}_{i}}=\mathcal{N}_{i}/G,$ and $\bar{f}_{i}:\mathcal{M}_{i}\to\mathcal{N}_{i}/G$ for the weak morphism induced by $f_{i}$, we would want a balanced map $\Psi:=\bar{f}_{1}\times\bar{f}_{2}:{\mathcal{M}_{1}}\times{\mathcal{M}_{2}}\to\overline{\mathcal{N}_{1}}\otimes\overline{\mathcal{N}_{2}}$ given by
@@
But $\Psi(v_{1}+w_{1},x_{2})=f_{1}(v_{1}+w_{1})Gf_{2}(x_{2})$ whereas $\Psi(v_{1},x_{2})\boxplus\Psi(w_{1},x_{2})=f_{1}(v_{1})Gf_{2}(x_{2})\boxplus f_{1}(w_{1})Gf_{2}(x_{2}),$ which could be larger. Thus, the tensor product of homomorphisms modulo the residue map is too naive in general (but see Corollary 5.9 below).
@@
In [20], to expedite a categorical approach for hypermagmas, the tensor product of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ is defined as satisfying
@@
If $v_{1}\boxplus v_{2}=v_{1}$ and $w_{1}\boxplus w_{2}=w_{1}$ then $((v_{1}\otimes w_{1})\boxplus(v_{1}\otimes w_{2}))\boxplus((v_{2}\otimes w_{1})\boxplus(v_{2}\otimes w_{2}))=(v_{1}\otimes w_{1})\boxplus(v_{2}\otimes w_{1})=(v_{1}\otimes w_{1})$ whereas
$(v_{1}\otimes w_{1})\boxplus((v_{1}\otimes w_{2})\boxplus(v_{2}\otimes w_{1}))\boxplus(v_{2}\otimes w_{2}))=(v_{1}\otimes w_{1})\boxplus(\emptyset\boxplus(v_{2}\otimes w_{2}))=\emptyset.$
Thus, although [20] is appropriate for $\preceq$-morphisms in the category of hypermagmas, we feel that the more traditional approach may be needed in the category of hypermodules.